2-A2 Cost-Volume-Profit and Vending Machines 


Vendmart Food Services Company operates and services snack vending machines located in restaurants, gas stations, and factories in four southwestern states. The machines are rented from the manufacturer. In addition, Vendmart must rent the space occupied by its machines. The following expense and revenue relationships pertain to a contemplated expansion program of 80 machines. 


Fixed monthly expenses follow: 

Machine rental: 80 machines @ $22.10  $1,768

Space rental: 80 locations @ $20.00 1,600

Part-time wages to service the additional 80 machines   500

Other fixed costs 132

Total monthly fixed costs   $4,000 


Other data follow: 


Per Unit (Snack) Per $100 of Sales   
Selling price$1.00  100% 
Cost of snack.6868
Contribution margin $.32 32% 


These questions relate to the given data unless otherwise noted. Consider each question independently.

 1. What is the monthly break-even point in number of units (snacks)? In dollar sales? 

2. If 45,000 units were sold, what would be the company's net income? 

3. If the space rental cost was doubled, what would be the monthly break-even point in number of units? In dollar sales? 

4. Refer to the original data. If, in addition to the fixed space rent, Vendmart Food Services Company paid the vending machine manufacturer $.07 per unit sold, what would be the monthly break-even point in number of units? In dollar sales? 

5. Refer to the original data. If, in addition to the fixed rent, Vendmart paid the machine manufacturer $.11 for each unit sold in excess of the break-even point, what would the new net income be if 45,000 units were sold?
Answer #1

1) Break even point units= Fixed cost/Contribution margin per unit

= $4000/0.32= 12500 snacks

Break even point in dollar sales= Fixed cost/Contribution margin ratio

= $4000/32%= $12500

2)

Sales revenue (45000*$1) $45000
Less: Variable costs (45000*0.68) 30600
Contribution margin 14400
Less: Fixed costs 4000
Net income $10400

3) Calculation of New fixed costs

Machine rental (80*$22.10) $1768
Space rental 80*($20*2) 3200
Part time wages 500
Other fixed costs 132
Total monthly fixed costs $5600

Break even point units= Fixed cost/Contribution margin per unit

= $5600/0.32= 17500 snacks

Break even point in dollar sales= Fixed cost/Contribution margin ratio

= $5600/32%= $17500

4) New variable costs= $0.68+0.07= $0.75

New contribution margin= Selling price-New variable costs

= $1.00-0.75= $0.25

Contribution margin ratio= Contribution margin*100/Sales

= $0.25*100/1= 25%

Break even point units= Fixed cost/Contribution margin per unit

= $4000/0.25= 16000 snacks

Break even point in dollar sales= Fixed cost/Contribution margin ratio

= $4000/25%= $16000

5) Additional variable cost in excess of break even point= (45000-12500)*0.11= $3575

New variable costs= $0.68*45000+3575= $34175

Contribution margin= Sales-New variable costs

= (45000*$1)-34175= $10825

Net income= Contribution margin-Fixed costs

= $10825-4000= $6825